13. Express -1+I in polar form. Hence show that(-1+i)¹⁶ is real and that 1/(-+i)⁶ is purely imaginary, giving the value for each
1
Expert's answer
2020-06-01T14:58:33-0400

a = r cos θ

b = r sin θ

a+bi= r (cos θ+i sin θ)

here a=-1 and b=1

here r="\sqrt{a^2+b^2}=\sqrt2"

θ="\tan^-(b/a)=\tan^-(-1)=3\pi/4"

so in polar form it becomes "\sqrt2(\cos\frac{3\pi}{4}+i\sin\frac{3\pi}{4})"


"((-1+i)^2)^8=(1-1-2i)^8=(-2)^8.(i)^8=2^8.1=2^8"

Hence it is purely real



Let's calculate the value of "\frac{1}{(-1+i)^6}"

"\frac{1}{(1+i^2-2i)^3}=\frac{1}{(-2i)^3}=\frac{1}{8i}=-\frac{i}{8}"

which is purely imaginary